OpenStudy (anonymous):

A projectile is thrown upward so that its distance above the ground after t seconds is h = -16t2 + 672t. After how many seconds does it reach its maximum height? 21 s 31.5 s 42 s 10 s

OpenStudy (superdavesuper):

OUCH Calculus :P OK so the height is at max when dh/dt = 0. Do u know how to differentiate h by t?

OpenStudy (anonymous):

Ouch indeed. I'm a bit clueless on this one. Should I be using a formula such as dv=4.9t^2 or dh=Vht?

OpenStudy (superdavesuper):

h = -16t2 + 672t so what is dh/dt =? com'on i know u can do it :)

OpenStudy (anonymous):

I came up with 42 seconds. Mind telling me if that's correct?

OpenStudy (superdavesuper):

Sorry...try the differentiation again plz? u are very very close @hales :)

OpenStudy (anonymous):

I'm completely perplexed. My best guess would be 21 seconds.

OpenStudy (superdavesuper):

BINGO :) Hahaa u got me concerned n thought u giving up lol dh/dt = -32t + 672 = 0 but u got it urself - good job!!

OpenStudy (anonymous):

Thanks so much! I appreciate the help and encouragement. (:

OpenStudy (agent0smith):

You don't need calculus for this, you might've learned how to do it in algebra 2... remember x= -b/(2a)? (it comes from calculus though)